OpenStudy (anonymous):

Cleo and Clare are looking from their balcony to a swimming pool below that is located 15 m horizontally from the bottom of their building. They estimate the balcony is 45 m high and wonder how fast they would have to jump horizontally to succeed in reaching the pool. What calculations would you show to help them determine the answer? Evaluate the practicality of their being able to succeed at jumping into the pool.

OpenStudy (jtfan2000):

s=ut+1/2at^2 45=0+1/2gt^2 45=4.9t^2 9.184=t^2 t = 3.03 seconds from balcony to splat